# Phase space analysis of CCDM cosmologies

### v1

@@ Abstract
We perform a detailed investigation of the CCDM (creation of cold dark matter) cosmologies using the powerful techniques of qualitative analysis of dynamical systems. Considering a wide variety of the creation rates ranging from constant to dynamical, we examine the nature of critical points and their stability obtained from the individual scenario consisting of only cold dark matter, or cold dark matter plus a second fluid with constant equation of state. According to our analyses, these scenarios predict unstable dark matter dominated critical points, stable accelerating attractors dominated either by dark matter or the second fluid, scaling attractors in which both dark matter and the second fluid co-exist. Along with these critical points, these scenarios also indicate the possibility of decelerating attractors or decelerating scaling attractors in the future which are new results in this direction. These altogether suggest that CCDM cosmologies are viable alternatives to the mainstream cosmological models.
@@ I Introduction
Accelerating expansion of the universe Riess et al. (1998); Perlmutter et al. (1999) is a major discovery of astrophysics and cosmology but the intrinsic mechanism behind this phenomenon is not clearly understood. Usually two well known approaches are considered in this regard: one is the introduction of some hypothetical dark energy (DE) fluid with sufficient negative pressure Caldwell et al. (1998); Caldwell (2002); Amendola (2000); Brax and Martin (2000); Astier (2001); Boyle et al. (2002); Kamenshchik et al. (2001); Linder (2003); Padmanabhan (2002); Debnath et al. (2004); Li (2004); Feng et al. (2006); Guo et al. (2005); Vikman (2005); Jassal et al. (2005); Gong (2004); Babichev et al. (2005); Chiba (2006); Cataldo et al. (2005); Zhang et al. (2006a, b); Zimdahl (2005); Scherrer (2006); Li et al. (2006); Guo et al. (2007); Cai (2007); Feng and Li (2009); Linder and Scherrer (2009); Li and Barrow (2009); Dutta et al. (2009); de Putter et al. (2010); Feng and Lu (2011); Ma and Zhang (2011); Solà and Gómez-Valent (2015); Das et al. (2018); Pan (2018); Pan et al. (2018); Yang et al. (2019); Almaraz et al. (2020); Yang et al. (2020); Saridakis (2020); Hernández-Almada et al. (2022); Yang et al. (2023); Kumar et al. (2023); Rezaei et al. (2024); Giarè et al. (2024) (also see Peebles and Ratra (2003); Copeland et al. (2006); Frieman et al. (2008); Li et al. (2011a, 2013a); Bamba et al. (2012a)) assuming that the Einstein’s General Relativity (GR) is the correct theory of gravity in the cosmological scales, and secondly, the modifications of the Einstein’s GR  Nojiri and Odintsov (2003); Nojiri et al. (2005); Nojiri and Odintsov (2005, 2006a); Amendola et al. (2007a); Li et al. (2007a); Brookfield et al. (2006); Nojiri and Odintsov (2007); Cognola et al. (2008); Amendola et al. (2007b); Li et al. (2007b); Tsujikawa (2008); Fay et al. (2007); Li et al. (2007c); de la Cruz-Dombriz et al. (2008); Brax et al. (2008); De Felice and Tsujikawa (2009); Thongkool et al. (2009); Zhou et al. (2009); Saridakis (2010); Miranda et al. (2009); Li et al. (2011b); Nojiri and Odintsov (2011); Li et al. (2011c); Paliathanasis et al. (2011); He et al. (2011); Geng et al. (2011); Li et al. (2011c); Harko et al. (2011); Li et al. (2013b); Chakraborty (2013); Bamba et al. (2012b); Odintsov and Sáez-Gómez (2013); He et al. (2013); Nojiri and Odintsov (2014); Thomas et al. (2015); Liu et al. (2016); Nunes et al. (2016); Paliathanasis et al. (2016); Nunes et al. (2017a); Kennedy et al. (2017); Nunes et al. (2017b); Hernández-Aguayo et al. (2019); Arnold et al. (2019); Atayde and Frusciante (2021); Lobato et al. (2021); Geng et al. (2022); Pan et al. (2021); dos Santos et al. (2022); Santos (2023); Davies et al. (2024) (known as Modified gravity, hereafter MG in short) can also explain this accelerating expansion (see the review articles in this direction Nojiri and Odintsov (2006b); De Felice and Tsujikawa (2010); Capozziello and De Laurentis (2011); Clifton et al. (2012); Koyama (2016); Cai et al. (2016); Nojiri et al. (2017); Akrami et al. (2021); Bahamonde et al. (2023)). Despite many efforts in constructing a variety of DE and MG models, an ultimate cosmological scenario consistent with all the observational datasets is still under the sea. Although the $\Lambda$-Cold Dark Matter ($\Lambda$CDM) model (constructed within the framework of GR) in which $\Lambda$ acts as DE, has been found to be quite successful with many astronomical probes, but the assumption of independent conservation of the cold dark matter (DM) and DE within this framework has no well founded explanation. It is well known that $\Lambda$CDM already faces many challenges, such as the cosmological constant problem Weinberg (1989), cosmic coincidence problem Zlatev and Steinhardt (1999), and cosmological tensions in recent times Di Valentino et al. (2021); Perivolaropoulos and Skara (2022); Abdalla et al. (2022). Thus, in principle, there is no reason to prefer any cosmological proposal over the other and new models are welcome provided they are found to be consistent with the observational data. An alternative scenario to both DE and MG approaches, namely, the theory of gravitationally induced adiabatic particle creation or matter creation was proposed in the literature Alcaniz and Lima (1999); Freaza et al. (2002); Steigman et al. (2009); Lima et al. (2010); Basilakos and Lima (2010); Lima et al. (2016) that can explain the late-time accelerating expansion of the universe. In fact, matter creation models can also explain the early inflationary era as well Abramo and Lima (1996); Gunzig et al. (1998) which is quite promising for a new cosmological scenario aiming to compete with existing cosmological proposals. The development of this theory started with a pioneering work by Prigogine, Geheniau, Gunzig and Nardone Prigogine et al. (1988) who established the link between the matter creation and the Einstein’s gravitational equations. [^1] We refer to Calvao et al. (1992); Lima and Germano (1992); Zimdahl and Pavon (1994); Gariel and le Denmat (1995); Lima et al. (1996); Lima and Abramo (1999) discussing the connection between thermodynamics and matter creation and some of its important aspects, such as the equivalence of matter creation and bulk viscosity Zimdahl (1996); Fabris et al. (2006); Colistete et al. (2007); Li and Barrow (2009). In matter creation theory, the creation pressure caused by the produced particles plays the central role in driving the accelerating expansion of the universe. The creation pressure is directly linked with the rate of particle creation and the properties of the matter component from which the particles are created. As a result of which, without adding any hypothetical fluid or modifying the underlying gravitational theory, one can explain the accelerating phase of the universe. While looking at the large amount of works in the direction of DE (Refs. Caldwell et al. (1998); Caldwell (2002); Amendola (2000); Brax and Martin (2000); Astier (2001); Boyle et al. (2002); Kamenshchik et al. (2001); Linder (2003); Padmanabhan (2002); Debnath et al. (2004); Li (2004); Feng et al. (2006); Guo et al. (2005); Vikman (2005); Jassal et al. (2005); Gong (2004); Babichev et al. (2005); Chiba (2006); Cataldo et al. (2005); Zhang et al. (2006a, b); Zimdahl (2005); Scherrer (2006); Li et al. (2006); Guo et al. (2007); Cai (2007); Feng and Li (2009); Linder and Scherrer (2009); Li and Barrow (2009); Dutta et al. (2009); de Putter et al. (2010); Feng and Lu (2011); Ma and Zhang (2011); Solà and Gómez-Valent (2015); Das et al. (2018); Pan (2018); Pan et al. (2018); Yang et al. (2019); Almaraz et al. (2020); Yang et al. (2020); Saridakis (2020); Hernández-Almada et al. (2022); Yang et al. (2023); Kumar et al. (2023); Rezaei et al. (2024); Giarè et al. (2024)), and MG (Refs. Nojiri and Odintsov (2003); Nojiri et al. (2005); Nojiri and Odintsov (2005, 2006a); Amendola et al. (2007a); Li et al. (2007a); Brookfield et al. (2006); Nojiri and Odintsov (2007); Cognola et al. (2008); Amendola et al. (2007b); Li et al. (2007b); Tsujikawa (2008); Fay et al. (2007); Li et al. (2007c); de la Cruz-Dombriz et al. (2008); Brax et al. (2008); De Felice and Tsujikawa (2009); Thongkool et al. (2009); Zhou et al. (2009); Saridakis (2010); Miranda et al. (2009); Li et al. (2011b); Nojiri and Odintsov (2011); Li et al. (2011c); Paliathanasis et al. (2011); He et al. (2011); Geng et al. (2011); Li et al. (2011c); Harko et al. (2011); Li et al. (2013b); Chakraborty (2013); Bamba et al. (2012b); Odintsov and Sáez-Gómez (2013); He et al. (2013); Nojiri and Odintsov (2014); Thomas et al. (2015); Liu et al. (2016); Nunes et al. (2016); Paliathanasis et al. (2016); Nunes et al. (2017a); Kennedy et al. (2017); Nunes et al. (2017b); Hernández-Aguayo et al. (2019); Arnold et al. (2019); Atayde and Frusciante (2021); Lobato et al. (2021); Geng et al. (2022); Pan et al. (2021); dos Santos et al. (2022); Santos (2023); Davies et al. (2024)), it is fairly understood that even though the DE and MG models got massive popularity in the community compared to the models arising from the theory of particle creation, however, the matter creation models are equally compelling due to their simplicity and elegant nature, see also de Roany and Freitas Pacheco (2011); Lima et al. (2011); Jesus et al. (2011); Lima et al. (2012, 2014); Lima and Baranov (2014); Ramos et al. (2014); Fabris et al. (2014); Chakraborty et al. (2014); Baranov and Lima (2015); de Haro and Pan (2016); Pan et al. (2016); Paliathanasis et al. (2017); Biswas et al. (2017); Bhattacharya et al. (2018); Pan et al. (2019); Ivanov and Prodanov (2019a, b); Singh and Kaur (2019); Cárdenas et al. (2020a, b); Gohar and Salzano (2021); Kaur and Singh (2021); Cárdenas et al. (2021, 2022); Pinto et al. (2023); Trevisani and Lima (2023); Banerjee et al. (2024); Montani et al. (2024); Elizalde et al. (2024); Mandal and Biswas (2025); Cárdenas and Lepe (2025).
The heart of this theory is the matter creation rate -- the rate at which matter particles are created. For a given particle creation rate, one can in principle determine the dynamics of the universe. However, one of the limitations of matter creation theory is the unavailability of any fundamental theory which can evaluate the particle creation rate, hence, different choices for the particle creation rate are considered and the resulting cosmological scenario is tested with the observational data. This treatment is almost identical with the DE and MG gravity theories where a hypothetical fluid (in the context of GR) or an unknown modified gravity is considered at the beginning, and no fundamental principle is available so far that can correctly describe everything. Even though, for some specific phenomenological choices of $\Gamma$, see for instance Lima et al. (2010, 2016), one can mimic the $\Lambda$CDM-like cosmology, however, based on the ongoing debates in the scientific community Di Valentino et al. (2021); Perivolaropoulos and Skara (2022); Schöneberg et al. (2022); Abdalla et al. (2022); Kamionkowski and Riess (2023), $\Lambda$CDM is not the final destination of the universe, and most probably, new physics beyond $\Lambda$CDM is needed. Therefore, examining different matter creation rates is equally compelling and worth exploring given the fact that currently, the dynamics of the universe is still elusive.
@@
The paper has been organized as follows. In section II, specifically in its first half, we describe the gravitational field equations for one fluid (section II.1) and two-fluid (section II.2) systems; and in the second half of this section, i.e. in section II.3 we present various matter creation models emerging from a general matter creation model. In section III we present the autonomous systems for the proposed matter creation models and examine the phase space analysis of the individual scenario. Finally, in section IV we conclude the present article with the main findings.
II Matter Creation Cosmology: One fluid and two-fluid systems
This section deals with the gravitational field equations of a matter creation cosmology. We consider two things to proceed with the matter creation cosmology. The first one is the validity of Einstein’s General Relativity in the cosmological scales and the second one is the homogeneous and isotropic universe which is well described by the spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) line element $ds^{2}=-{dt}^{2}+a^{2}(t)\left[dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\right]$ in which $a(t)$ is the expansion scale factor of the universe. Now, let us assume a system having a volume $V$ attaining $N$ particles. In an open thermodynamical system, $N$ should be a function of time i.e. it changes with time. Thus, the conservation equation will transform to the form
@@ II Matter Creation Cosmology: One fluid and two-fluid systems
where $n=N/V$ is the particle density, $N^{\mu}=nu^{\mu}$ corresponds to the particle flow vector ($u^{\mu}$ represents the four velocity of the particle); ‘dot’ stands for the derivative with respect to the cosmic time $t$; $\Theta=u^{\mu}_{;\mu}$ is the expansion scalar of the fluid which in the FLRW universe becomes $3H$  ($H=\dot{a}(t)/a(t)$ is the Hubble rate of the FLRW universe), and $\Gamma$ is the rate of change of particle number. Here $\Gamma>0$ indicates the production of particles while $\Gamma<0$ indicates the reverse, that means particle annihilation. Now, using (1) in the Gibb’s equation Zimdahl (1996)
@@
Considering an adiabatic (also known as isentropic) thermodynamic process where the rate of change of specific entropy vanishes (i.e. $\dot{s}=0$), eqn. (3) reduces to
@@
where the new term $p_{c}$ is termed as creation pressure due to the particle production and it takes the form
$p_{c}=-\frac{\Gamma}{3H}(\rho+p).$ (6)
This creation pressure has some interesting implications: if the fluid under consideration is a normal fluid, then for $\Gamma>0$, one can induce a negative creation pressure. As we shall explain in the following, this creation pressure can drive the accelerating phase of the universe and hence particle production scenario can be treated as a mirage of a DE. Now, concerning the matter sector, one can assume that i) the matter sector is comprised of only one fluid, for instance the pressure-less dark matter (DM), and this DM is endowed with gravitationally induced adiabatic matter creation Lima et al. (2010), or ii) the matter sector of the universe is comprised of $n$ arbitrary fluids and all of them can create gravitationally induced adiabatic matter particles Lima et al. (2016), or alternatively, ii) the matter sector is comprised of two mixed fluids where some fluids are able to create gravitationally induced adiabatic matter particles while the remaining fluids do not take part in the matter creation. Theoretically, all of the above possibilities are equally valid and appealing. In this work we shall consider two different matter creation cosmologies, namely, one fluid system in which the matter sector of the universe is comprised of only pressure-less DM sector blessed with gravitationally induced adiabatic matter creation, and secondly we shall consider a two-fluid system made of a pressure-less DM (responsible for the gravitationally induced adiabatic matter creation) and a general fluid with a barotropic equation of state which does not take part in the matter creation hypothesis; and in addition to that, both the fluids do not take part in an energy exchange mechanism, that means they are not interacting with each other. The introduction of the second fluid is motivated from the fact that the phase space structure of the matter creation cosmologies might be affected in presence of a second fluid which does not join in the matter creation process.
II.1 One fluid system
For the single fluid system, namely, pressure-less DM endowed with gravitationally induced adiabatic particle production, in the context of the FLRW background, the gravitational field equations become,
$H^{2}=\frac{\kappa^{2}}{3}\;\rho_{\rm dm},$ (7)
$2\dot{H}+3H^{2}=-{\kappa}^{2}p_{c},$ (8)
where $\kappa^{2}=8\pi G$ is the Einstein’s gravitational constant ($G$ is the Newton’s gravitational constant); $\rho_{\rm dm}$ is the energy density of the pressure-less DM, $p_{c}$ is the creation pressure which is related to the rate of particle production $\Gamma$ as Steigman et al. (2009); Lima et al. (2010); Basilakos and Lima (2010); de Haro and Pan (2016)
@@ II.1 One fluid system
Here $\Gamma>0$ indicates the creation of particles and $\Gamma<0$ indicates the particle annihilation. [^2] Since the pressure-less DM sector is only responsible for the creation, so it is natural to depict that the created particles are the pressure-less DM particles. It is interesting to note that the conservation equation of the pressure-less DM which follows from (7) and (8) is $\dot{\rho}_{\rm dm}+3H\left(1-\Gamma/3H\right)\rho_{\rm dm}=0$, is equivalent to the conservation equation of a non-cold DM fluid with variable equation of state of DM: $\dot{\rho}_{\rm dm}+3H(1+w_{\rm dm}(t))\rho_{\rm dm}$, where $w_{\rm dm}(t)=-\Gamma/3H$. Thus, the particle creation mechanism can be considered to be an equivalent prescription of the non-interacting cosmological scenario with a variable equation of state of the underlying fluid (here DM) responsible for this creation. Now, using (7), (8), and (9) one arrives at the acceleration equation
$\frac{\ddot{a}}{a}=-\frac{\kappa^{2}}{6}\left(1-\frac{\Gamma}{H}\right)\rho_{\rm dm},$ (10)
from which one can see that if $\Gamma<0$, then $\ddot{a}<0$ representing a decelerating phase of the universe. In order to realize an accelerating expansion of the universe, $\Gamma$ needs to satisfy $01-\Gamma/H<$ 0, which indicates that in an expanding universe (i.e. $H>0$), $\Gamma>H$. That means, the rate of particle creation should be positive.
II.2 Two-fluid system
In the two-fluid system, as already mentioned, one fluid is the pressure-less DM which is gifted with the gravitationally induced adiabatic matter creation and the second fluid does not take part in this matter creation process and both of them are independently conserved. The gravitational equations in the context of the FLRW universe become,
$H^{2}=\frac{\kappa^{2}}{3}(\rho_{\rm dm}+\rho_{f}),$ (11)
$2\dot{H}+3H^{2}=-{\kappa}^{2}(p_{c}+p_{f}),$ (12)
@@ II.2 Two-fluid system
where $\rho_{f}$, $p_{f}$ are respectively the energy density and pressure of the second fluid; the creation pressure $p_{c}$ associated with the DM fluid is related to the particle creation or matter creation rate $\Gamma$ as in (9). The conservation equation of both the fluids become, $\dot{\rho}_{\rm dm}+3H(\rho_{\rm dm}+p_{c})=0$, and $\dot{\rho_{f}}+3H(1+w)\rho_{f}=0$, where $w=p_{f}/\rho_{f}$ is the barotropic equation-of-state of the second fluid and it could take any real number. Let us note that similar to the one-fluid system, the matter creating DM fluid is equivalent to a non-interacting DM fluid with variable equation of state. Now, concerning $w$, if it is a constant, then one can find the evolution of $\rho_{f}$ as, $\rho_{f}=\rho_{f0}(a/a_{0})^{-3(1+w)}$ where $\rho_{f0}$ is the present value of $\rho_{f}$ and $a_{0}$ is the scale factor at present time. If on the contrary, $w$ is dynamical, then depending on the nature of $w$, one can either solve $\rho_{f}$ analytically or numerically. However, the overall evolution of this matter creating cosmology depends on the matter creation rate $\Gamma$. In this work we focus on the constant $w$ in order to start with a simple scenario of matter creation cosmology. Finally, the accelerating equation in this case becomes,
$\frac{\ddot{a}}{{a}}=-\frac{\kappa^{2}}{6}\left[\left(1-\frac{\Gamma}{H}\right)\rho_{\rm dm}+(1+3w)\rho_{f}\right].$ (13)
@@
From eqn. (13) one can clearly see that if we consider a normal fluid characterized by $w\geq 0$, then for $\Gamma<0$, we will never realize an accelerating expansion of the universe because in this case $\ddot{a}<0$, however, for $\Gamma>0$, it is possible to get $\ddot{a}>0$ provided the sum of the terms inside the third brace of the right hand side of eqn. (13) becomes negative. This can be obtained for some suitable choices of $\Gamma$. However, we further note that for $\Gamma<0$ it is not impossible to obtain the accelerating expansion of the universe, because in this case we need to allow $w<0$, that means particle annihilation alone cannot explain the accelerating expansion unless we add some hypothetical fluid from outside. This further strengthens the presence of a hypothetical fluid with negative pressure in the context of particle annihilation. While the focus of this work is not to consider some hypothetical fluid with negative pressure (i.e. $w<0$); we are mainly interested to investigate the effects of $\Gamma~{}(>0)$ on the accelerating expansion of the universe in presence of a normal fluid (i.e. $w\geq 0$). However, in the present article we shall allow both the non-negative and negative values of $w$ in order to compare the effects of $\Gamma$ on the resulting scenarios.
II.3 Models
Over the last couple of years, a variety of matter creation rates have been proposed in the literature Lima et al. (2011); Jesus et al. (2011); Lima et al. (2012, 2014); Lima and Baranov (2014); Chakraborty et al. (2014); Baranov and Lima (2015); de Haro and Pan (2016); Pan et al. (2016); Paliathanasis et al. (2017); Pan et al. (2019). In most of the cases, the choice of the matter creation rate $\Gamma$ is arbitrary since there is no fundamental guiding principle available yet in the literature, however, the resulting phenomenological scenarios become very appealing. A general choice of the matter creation follows
$\Gamma=\sum_{j=1}^{\infty}\Gamma_{-j}(\mathcal{X})H^{-j}+\Gamma_{0}+\sum_{i=1}^{\infty}\Gamma_{i}(\mathcal{X})H^{i}$ (14)
where $\mathcal{X}$ could be any cosmic variable, $\Gamma_{-j}(\mathcal{X})$’s, $\Gamma_{j}(\mathcal{X})$’s may not necessarily be dimensionless. We can only say that $\Gamma$ has the dimension of the Hubble rate. Thus, depending on the presence of $H^{i}$’s in the expression (14), the dimension of $\Gamma_{i}(\mathcal{X})$’s will be decided consequently. In order to proceed with the phase space analysis, a typical matter creation rate needs to be provided. Although one can start with a general matter creation rate from (14) which may involve the arbitrary powers of $H$, however, at the same time, the number of parameters will increase and understanding the phase space structure might be complicated. Thus, in the first half of the article we consider some simple choices of the matter creation rate as follows,
$\Gamma=\Gamma_{0},$ (15)
$\Gamma=\Gamma_{0}H,\quad\Gamma=\Gamma_{0}H^{2},$ (16)
$\Gamma=\Gamma_{0}H^{-1},\quad\Gamma=\Gamma_{0}H^{-2},$ (17)
and explain how different choices of the matter creation rate affects the phase space structure. In the second half of the article we consider various matter creation rates which are the linear combinations of $H$, $1/H$, $H^{2}$ and investigate the phase space analysis. We consider the following choices of the matter creation rate:
$\Gamma=\Gamma_{0}+\Gamma_{1}H,$ (18)
$\Gamma=\Gamma_{0}+\Gamma_{3}H^{-1},$ (19)
$\Gamma=\Gamma_{0}+\Gamma_{2}H^{2},$ (20)
$\Gamma=\Gamma_{1}H+\Gamma_{2}H^{2},$ (21)
$\Gamma=\Gamma_{1}H+\Gamma_{3}H^{-1},$ (22)
$\Gamma=\Gamma_{0}+\Gamma_{1}H+\Gamma_{2}H^{2}.$ (23)
With these matter creation models, one can cover a wide variety of models in the literature. As far as we are concerned with the literature, only a few models have been considered in the literature. Having these matter creation models, in the following we perform the phase space analysis of the individual model.

### v2

# Matter Creation Cosmologies and Accelerated Expansion

@@ Abstract
Non-conservation of dark matter can lead to late-time cosmic acceleration. This mechanism is known as the matter creation theory and this replaces the need of dark energy and modified gravity theories. We consider a two-fluid system consisting of a cold dark matter and a second fluid with constant barotropic equation of state. We performed detailed investigations of such cosmologies using the powerful techniques of qualitative analysis of dynamical systems. Considering a wide variety of the creation rates, we examine the phase space analysis of the individual scenario. According to our analyses, these scenarios predict decelerating unstable dark matter (or second fluid) dominated critical points, accelerating attractors dominated either by dark matter or the second fluid, accelerating scaling attractors in which dark matter and the second fluid co-exist. The regime of late-time accelerating expansion can be classified as either quintessence, phantom or driven by a cosmological constant. This huge variety of critical points makes these scenarios phenomenologically rich, and naturally suggests that such scenarios can be viewed as viable and potential alternatives to the mainstream cosmological models.
Cosmology; Matter creation; Dark matter; Dynamical system analysis; Critical points
@@ I Introduction
[…]
@@
The paper has been organized as follows. In section II, we describe the gravitational equations of a two-fluid matter creation scenario and present the matter creation models that we wish to study. In section III we present the autonomous systems for the proposed matter creation models and examine the phase space analysis of the individual scenario. Finally, in section IV we conclude the present article with the main findings.
II Matter Creation Cosmology
We assume that the gravitational sector of the universe is described by the Einstein’s General Relativity and our universe in its large scale is described by the Friedmann-Lemaître-Robertson-Walker (FLRW) line element
$ds^{2}=-{dt}^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\right],$ (1)
where $a(t)$ is the expansion scale factor of the universe and $(t,r,\theta,\phi)$ are the co-moving coordinates. Here $k$ is the spatial curvature and its null value corresponds to the spatially flat universe while k=1k=-1 indicates that our spatial geometry of the universe is open while $k=+1$ corresponds to the closed spatial geometry of the universe. We consider a system having a volume $V$ which attain $N$ number of particles. In an open thermodynamical system, $N$ should be a function of time, i.e. it should change with time. Thus, the conservation equation will transform to the form
@@ II Matter Creation Cosmology
where $n=N/V$ is the particle density, $N^{\mu}=nu^{\mu}$ corresponds to the particle flow vector ($u^{\mu}$ represents the four velocity of the particle); ‘dot’ stands for the derivative with respect to the cosmic time $t$; $\Theta=u^{\mu}_{;\mu}$ is the expansion scalar of the fluid which in the FLRW universe becomes $3H$  ($H=\dot{a}(t)/a(t)$ is the Hubble rate of the FLRW universe), and $\Gamma$ is the rate of change of particle number. Here $\Gamma>0$ indicates the production of particles while $\Gamma<0$ indicates particle annihilation. Using eqn. (2) in the Gibb’s eqn. [134]
@@
Under the adiabatic (also known as isentropic) thermodynamic process where the rate of change of specific entropy vanishes (i.e. $\dot{s}=0$), eqn. (4) reduces to
@@
where the new term $p_{c}$ is termed as the creation pressure due to the particle production and it takes the form $p_{c}=-(\Gamma/3H)\times(p+\rho)$. This creation pressure has some interesting implications: if the fluid under consideration is a normal fluid, then for $\Gamma>0$, one can induce a negative creation pressure. As we shall explain in the following, this creation pressure can drive the accelerating phase of the universe and hence particle production scenario can be treated as a mirage of a DE. Now, we assume that the total energy density of the universe is comprised of a pressure-less DM sector endowed with gravitationally induced adiabatic matter creation, and a second fluid which does not take part in the matter creation hypothesis. Additionally we consider that the fluids under consideration do not take part in an energy exchange mechanism, that means they are not interacting with each other. The gravitational equations in the context of the FLRW universe now explicitly be written as
$H^{2}+\frac{k}{a^{2}}=\frac{\kappa^{2}}{3}(\rho_{\rm dm}+\rho_{f}),$ (7)
$2\dot{H}+3H^{2}+\frac{k}{a^{2}}=-{\kappa}^{2}(p_{c}+p_{f}),$ (8)
where $\kappa^{2}=8\pi G$ is the Einstein’s gravitational constant ($G$ is the Newton’s gravitational constant); $\rho_{f}$, $p_{f}$ are respectively the energy density and pressure of the second fluid; $\rho_{\rm dm}$ is the energy density of the pressure-less DM, $p_{c}$ is the creation pressure which is related to the rate of particle production $\Gamma$ as [121, 122, 123, 147]
@@
Here $\Gamma>0$ indicates the creation of particles and $\Gamma<0$ indicates the particle annihilation. [^2] Since the DM sector is the only responsible fluid for the creation, it is natural to depict that the created particles are the DM particles. The conservation equation of the second fluid is, $\dot{\rho_{f}}+3H(1+w)\rho_{f}=0$, where $w=p_{f}/\rho_{f}$ is the barotropic equation-of-state (EoS) of the second fluid and here we assume it to be a constant. This EoS is indefinite in sign. Although this fluid does not take part in the matter creation but it influences the expansion history of the universe, therefore, it is evident that the EoS, $w$, plays a crucial role in this context. In particular, the dynamics of the universe in presence of a normal fluid (characterized by $w\geq 0$) and hypothetical fluid (characterized by $w<0$) should be different. Considering this, in this article we explore all kind of possibilities through its EoS as follows: $0\leq w\leq 1$ (this corresponds to a normal fluid where $w=1$, $w=1/3$, $w=0$ respectively corresponds to a stiff fluid, radiation and dust), $1/3\leq w<0$ (this is a hypothetical fluid in the sense that the pressure is negative), $1<w\leq-1/3$ (quintessence-like fluid), w=1w=-1 (it corresponds to the cosmological constant), w<1w<-1 (phantom fluid). While our main intention is to focus on the matter creation scenarios in presence of a normal fluid, however, we consider the remaining cases for completeness. On the other hand, it is interesting to note that the matter creating DM fluid is equivalent to a non-interacting DM fluid with variable EoS. This can be quickly understood by recasting the conservation equation of the DM fluid as
$\dot{\rho}_{\rm dm}^{\rm eff}+3H(1+w_{\rm dm}^{\rm eff})\rho_{\rm dm}^{\rm eff}=0,$ (10)
@@
where $w_{\rm dm}^{\rm eff}=-\Gamma/3H$ represents the effective EoS for DM and considering this effective description of DM, we use $\rho_{\rm dm}^{\rm eff}$ to represent $\rho_{\rm dm}$ and from now on we shall use $\rho_{\rm dm}^{\rm eff}$. For $\Gamma>0$, $w_{\rm dm}^{\rm eff}$ becomes negative in an expanding universe. This is not surprising since the possibility of negative EoS of DM has recently been reported in several articles [171, 172, 173, 174, 175, 176]. This establishes the equivalency of matter creation cosmologies and DM system with dynamical EoS. Now, concerning the second fluid, one can find the evolution of $\rho_{f}$ as, $\rho_{f}=\rho_{f0}(a/a_{0})^{-3(1+w)}$ where $\rho_{f0}$ is the present value of $\rho_{f}$ and $a_{0}$ is the scale factor at present time which without any loss of generality, we set $a_{0}=1$. If, on the contrary, $w$ is dynamical, then depending on the nature of $w$, one can either solve $\rho_{f}$ analytically or numerically. However, the overall evolution of this matter creating cosmology depends on the matter creation rate $\Gamma$. In this work we focus on the constant $w$ in order to start with a simple scenario of matter creation cosmology. Finally, the accelerating equation in this case becomes,
$\frac{\ddot{a}}{{a}}=-\frac{\kappa^{2}}{6}\left[\left(1-\frac{\Gamma}{H}\right)\rho_{\rm dm}^{\rm eff}+(1+3w)\rho_{f}\right].$ (11)
@@
From eqn. (11) one can clearly see that if we consider a normal fluid characterized by $w\geq 0$, then for $\Gamma<0$, we will never realize an accelerating expansion of the universe because in this case $\ddot{a}<0$; however, for $\Gamma>0$, it is possible to get $\ddot{a}>0$ provided the sum of the terms inside the third brace of the right hand side of eqn. (11) becomes negative. This can be obtained for some suitable choices of $\Gamma$. However, we further note that for $\Gamma<0$, it is not impossible to obtain the accelerating expansion of the universe, because in this case we need to allow $w<0$, that means particle annihilation together with some hypothetical fluid with negative pressure can drive the cosmic acceleration. This further strengthens the presence of a hypothetical fluid with negative pressure in the context of particle annihilation. However, in the present article we restrict ourselves to $\Gamma>0$ and explore the cosmological scenarios for both $w\geq 0$ and $w<0$.
[…]
Usually one can consider a number of choices for $\Gamma$ and investigate the viability of the resulting scenario. In the present work we consider the following two choices of $\Gamma$:
$\Gamma=\Gamma(H),\quad\quad\quad\Gamma=\gamma_{H}\dot{H},$ (13)
[…]
$\rho_{\rm dm}^{\rm eff}=\rho_{\rm dm,0}^{\rm eff}a^{-3}\exp\left(\int_{1}^{a}\frac{\Gamma}{aH}\,da\right)$ (14)
where $\rho_{\rm dm,0}^{\rm eff}$ is the present-day value of the DM density. While for the remaining matter creation rate, one can solve the energy density for DM as
$\rho_{\rm dm}^{\rm eff}=\rho_{\rm dm,0}^{\rm eff}a^{-3}\exp\Bigl[\gamma(H)-\gamma(H_{0})\Bigr]$ (15)
It is interesting to mention that with the proper choices of $\Gamma$, known cosmological scenarios can be recovered. Concerning the variable $\gamma(H)$ we consider two choices [156]:
$\gamma(H)=\left(1-\frac{1}{\xi}\right)\ln\left(\frac{3H^{2}}{H_{0}^{2}}\right),$ (16)
where $\xi\neq 0$ and
[…]
As noticed in Ref. [156], the above two choices of $\gamma(H)$ can lead to some interesting consequences. In the following we perform the phase space analysis of the individual model assuming the spatially flat FLRW universe (i.e. $k=0$).